OpenStudy (anonymous):

Hi i need solution of this integral ,,, Thank you . g(x)=\int\limits_{}^{}(-e ^{\frac{ x ^{2} }{2 }}\cos(x) )dx

OpenStudy (anonymous):

\[g(x)=\int\limits\limits_{}^{}(-e ^{\frac{ x ^{2} }{2 }}\cos(x) )dx\]

OpenStudy (dumbcow):

you cant integrate e^x^2 by hand, only using numerical methods http://www.wolframalpha.com/input/?i=integrate+-e%5E(x%5E2%2F2)*cos(x)+dx

OpenStudy (anonymous):

my friend i know this website , and i know the result already ,,, i know its hard to integral this ,, because the prof. challenge us to do it ,, also he said its difficult . i have an idea of making cos(x) in term of extensional .... its seems to be work ,,,,,,,Thanks for your help @dumbcow

OpenStudy (anonymous):

\[\int\limits_{}^{}\ (-e ^{\left( \frac{ x ^{2} }{2 } \right)} \left( \frac{ e ^{2i}+e ^{-2i} }{2} \right))dx=(-\frac{ 1 }{2 }) \int\limits_{}^{}(e^{x^2+2i} +e^{x^2-2i})dx=\left( \frac{ -1 }{2 } \right)\left( \int\limits_{}^{} e^{x^2+2i}dx+\int\limits e^{x^2-2i} dx\right)\]

OpenStudy (anonymous):

i believe its not need iteration method or whatever ,,, thank you everybody :)